# “When to Hand Off, When to Work Together”: Expanding Human-Agent Co-Creative Collaboration through Concurrent Interaction

### 06 Mar 2026 v2

@@ 3.2. Study Design
Our first study aims to investigate two key questions: (1) How does process transparency—observing the agent’s step-by-step workflow on the canvas—affect designers’ collaboration experience and enable a new type of collaboration? (2) What design considerations are needed to support better human-agent collaboration in co-creative design tasks?
To answer these questions, we recruited 10 professional UI/UX designers and had them work with our design probe on two types of design tasks, each representing a different level of task granularity. Through task observations and semi-structured interviews, we investigated how process transparency affects designers’ understanding, collaboration strategies, and expectations for agent behavior.
@@ 3.2.1. Participants
We recruited 10 UI/UX designers (D1-D10, 3 male, 7 female, ages 24-36, mean=28.30) with professional design experience. All participants had at least 1 year of professional experience or had completed at least three web or UI/UX design projects. We screened these participants for high Figma proficiency (e.g., daily users) and prior experience with diverse design agents (e.g., Figma Make (Make, ) or Lovart (Lovart, )). Participants were compensated 40,000 KRW for approximately 1.5 hours of participation.
@@ 3.2.2. Procedure
After introducing different types of design agents (e.g., OptiMuse (Zhou et al., 2024) and Figma Make (Make, )) and providing a brief tutorial on the design probe’s three interaction components (voice instruction, process observation, and presence visualization), participants completed two design tasks of varying scope: (1) a fine-grained task, which involved designing a single webpage section (e.g., hero or project gallery section) with given design materials (15 minutes), and (2) a coarse-grained task, which involved creating an entire webpage (e.g., startup landing page) including multiple sections (25 minutes). These tasks were designed to observe how collaboration methods differ between single-section and multi-section work. We concluded with a 30-minute semi-structured interview exploring how observing the agent’s workflow affected their collaboration strategies, their experience instructing the agent, and desired improvements for the current design probe.
$3.$ Findings
@@ 3.3.1. Process Transparency: Real-Time Understanding and Intervention Opportunities
Observing the agent’s step-by-step execution fundamentally transformed how designers experienced collaboration. All designers reported a strong sense of working with the agent rather than delegating to it. Unlike traditional agents, where users assign tasks and wait for final outputs, process transparency enabled simultaneous work: participants understood what the agent was doing and planned their own tasks accordingly, maintaining active engagement throughout execution.
This transparency relieved the “loss of agency” problem common in existing agents (Feng et al., 2024). Participants contrasted the first probe system with general design agents, where agents take over the workspace, leaving users waiting passively and fearing disruption if they intervene. With those agents, users must read lengthy reasoning texts to understand its plans—a cognitively demanding process. In contrast, seeing intermediate visual actions on the canvas made the agent’s intentions immediately understandable with minimal effort. D4 commented “I don’t need to read paragraphs—I just see it creating a button, and I know what’s happening.” This visual feedback preserved user agency by avoiding both over-dependence on opaque automation and wholesale rejection of agent outputs.
Most critically, step-by-step observation enabled early evaluation and proactive intervention. Participants could anticipate the agent’s next steps, identify the agent’s misinterpretation on the given task (D1, D3, and D5-D7), recognize their own unclear instructions (D1-D2, D4, and D8), and assess the agent’s overall workflow direction (D2, D6, and D8-D10)—all before the agent finished. Importantly, this understanding enabled participants to plan their own parallel work: while the agent executed, participants gathered reference images, evaluated overall design directions, or prepared content for subsequent sections, rather than passively waiting. Specifically, D7 explained: “I saw it working on the hero section with a certain direction, so I started looking for references related to that direction for the next section instead of just watching.” This predictability helped participants manage workflows more effectively, deciding when to intervene, when to prepare subsequent tasks, and when to let the agent continue autonomously. Process transparency thus shifted collaboration from reactive correction of final outputs to proactive coordination during execution.
@@ 3.3.2. Emerging Expectations: Flexible Concurrent Collaboration and Context Awareness
Agent’s process transparency fundamentally changed designers’ expectations: they wanted to engage directly with the agent’s work-in-progress through concurrent intervention rather than waiting for completion. For example, participants wanted to appropriate intermediate outputs for their own designs, assist the agent by working on related elements simultaneously, or demonstrate desired changes through direct editing. Most of the participants (D1-D3 and D6-D10) explained that observing the process made them more aware of how they wanted things done, creating natural moments to intervene and collaborate. D3 noted: “I could see exactly where it needed my input—I wanted to just jump in and show it rather than wait and explain later.”
However, the first probe’s inability to distinguish user actions from agent actions created conflicts. Because the agent tracked overall canvas changes without separating user modifications, it sometimes reverted or misinterpreted concurrent user work. Participants who experienced this avoided intervening during execution, relegating their collaboration ideas to post-task interviews. Yet even participants who avoided intervention reported that their reluctance to interrupt agents stems from prior experiences with other agent systems (D2, D4-D7, D10)—fearing that mid-execution changes would derail the entire agent’s workflow. Process transparency challenged this assumption; designers recognized opportune moments for intervention but lacked system support.
Participants envisioned real-time interleaved workflows where users could intervene and interleave their work with the agent’s ongoing execution: appropriating agent-created elements into their own designs while the agent continued working, handing off their work-in-progress for the agent to build upon mid-execution, or collaboratively editing the same artifact simultaneously. Critically, they expected agents to understand why they intervened—whether corrections indicated misunderstanding, appropriations showed approval, or demonstrations communicated desired style (D1, D3-D5, D7-D8). This requires collaborative context awareness, which means distinguishing feedback on agent work from independent user work and recognizing when users want the agent to adapt versus continue autonomously. Process transparency thus revealed not just new collaboration opportunities, but the need for context-aware agents that support—rather than conflict with—real-time interleaved human actions.
@@ 4. Study 2: Two-Day Technology Probe Study with Stimulated Recall Interviews
In this section, we describe an exploratory study that investigates when designers delegate hands-off tasks to agents, work collaboratively alongside them, or provide only verbal feedback during co-creative collaboration. Through a two-day technology probe study (Hutchinson et al., 2003) incorporating stimulated recall interviews (Dempsey, 2010; Calderhead, 1981) with creative professionals having over two years of experience, we sought to understand how designers navigate various action patterns in collaborative contexts with the agent, where natural, simultaneous work with agents is possible. This section first describes the refined design probe used in this study, followed by the study design, data analysis methods, and findings.
4.1. Second Probe System: Cleo (Collaborative Linked Executive Operator)
@@ 4.1.1. Understanding Collaborative Context for Flexible Human-Agent Co-Creation
While our first design probe includes three components (voice modality, process transparency, and presence awareness), supporting real-time interleaved workflows requires agents to understand and adapt to diverse human interventions that occur during the agent’s execution. Based on Study 1 findings, users expect to interleave their actions with the agent’s ongoing execution rather than waiting for task completion when they can observe and understand the agent execution process. The key challenge is collaborative context awareness—agents must distinguish whether concurrent user actions represent feedback on the agent’s work requiring adaptation, the user’s independent parallel work that should be preserved, or collaborative co-editing that should be incorporated into the agent’s ongoing execution. This requires that the agent should understand which elements users modify, whether they build upon or diverge from the agent’s work, and how users’ demonstrations relate to the ongoing task.
Therefore, we developed a second design probe, Cleo (Collaborative Linked Executive Operator), which extends the first probe by incorporating collaborative context awareness as its core principle. This enables the agent to recognize the user’s direct manipulation of the agent’s execution process and adapt its behavior accordingly while maintaining coherent shared progress.
@@ 4.1.2. Cleo Design and Improvements from Design Probe 1
Figure 3. Flexible co-creation scenario with Cleo. (a) The user invokes the agent by calling its name. Cleo allows users to (b) abort the agent’s operation, (c) work concurrently with the agent, or (d) intervene directly with additional instructions at any time.
@@
Interaction Improvements
Building on the first probe system, Cleo enables more dynamic and flexible collaboration as illustrated in Figure 3. Users can invoke the agent by saying “Cleo” without clicking the microphone button. The system then detects voice input and tracks canvas selections until several seconds of silence occur.
Cleo supports three types of interactions. First, users can abort the agent’s operation at any time via the termination button in the plugin interface (Figure 3B). Second, users can work concurrently with the agent, even on the same elements (Figure 3C). The agent tracks and prioritizes user operations over its own actions, updating its plan to align with user modifications. For instance, as shown in Figure 3C(2)-(3), when a user changes the title color to blue while the agent works on the menu, the agent incorporates blue as a key color in subsequent actions, such as adding blue decorative lines. Third, users can provide additional voice instructions at any time by calling the agent’s name (Figure 3D), which the agent incorporates in the next iteration.
Technical Improvements
To support these additional interactions, we updated the agent pipeline as shown in Figure 4. Instead of the original Summary Module, which only considers the agent’s tool calls and results, we introduced three modules to incorporate user actions into the agent’s workflow planning: Attribution Change Module, User Change Detection Module, and Plan Update Module.
When the agent receives user input, the User Change Detection Module analyzes user actions that occurred between the previous agent response and the current canvas state, capturing changes made while the agent was inactive. This allows the agent to maintain awareness of the overall workflow. The Attribution Change Module distinguishes between agent operations and user modifications by analyzing the canvas context after tool execution. It produces two separate summaries—one for agent actions and one for user actions during the iteration. These summaries are passed to the Plan Update Module, which updates the current plan only when user modifications conflict with the existing plan or indicate new directions, before initiating the next iteration cycle.
Figure 4. Agent structure of Cleo. Three modules have been updated from the first probe pipeline: User Change Detection Module, Attribution Change Module, and Plan Update Module.
@@ 4.2. Study Design
We conducted a two-day exploratory study to observe how creative professionals choose interaction approaches—whether to work together, fully delegate, or observe—when using Cleo. Based on these observations, we aimed to identify users’ action patterns during agent execution and understand why they chose each action. Specifically, by collecting interaction logs with Cleo and conducting stimulated recall interviews on the second day based on these logs, we aimed to investigate in detail the situations and rationales that led to particular interaction patterns. This understanding will enable future agent systems to recognize these patterns and adapt their execution strategies accordingly—knowing when to seek guidance, when to work autonomously, and when to support concurrent collaboration.
On the first day, participants performed two different types (e.g., closed- and open-ended) of web design tasks in Figma for approximately 70 minutes total. During this task session, we logged and screen-recorded the user’s and agent’s interactions. After this session, three authors visually represented the interaction logs in an easily reviewable format (Figure 5C-D), organizing the data into units. We introduced a one-day gap between the task session and recall interview for two reasons: first, to avoid participant fatigue from an extended single session (which would exceed three hours), and second, because preparing the visualized interaction logs required substantial processing time—the recorded video itself did not clearly reveal interaction dynamics between user and agent actions (especially what the agent was doing), necessitating manual timeline construction to show when the agent was executing specific operations and what the user was doing simultaneously. On the second day, stimulated recall interviews were conducted where participants reviewed these visualized interaction logs.
@@ 4.2.1. Participants
We recruited participants who met two key criteria: (1) at least two years of professional design experience, and (2) active use of VLMs (e.g., ChatGPT  [^6] https://chatgpt.com/ or Gemini  [^7] https://gemini.google.com/app) or design agents (e.g., Figma Make (Make, ), Stitch (Stitch, ), or Lovable (Lovable, )) in their creative workflow. A total of 10 participants were recruited (4 males and 6 females; M_age = 30.1, SD_age = 3.9), with an average of 3.8 years of professional experience (SD = 1.69, MIN = 2, MAX = 7). All participants were screened to ensure they had prior experience using not only general AI, such as ChatGPT, but also Figma Make (Make, )–specific AI design tools. Participants were compensated 120,000 KRW for approximately 3 hours of participation.
@@ 4.2.2. Procedure
First day
On the first day, after a 10-minute introduction, participants spent 30 minutes familiarizing themselves with the system through a warm-up task where they redesigned their personal web portfolio or company homepage using the design probe. Following this, the participants performed two main tasks. The first task was a 30-minute closed-ended design task where participants were given design requirements and source materials to design an empty section of a web portfolio to match other existing sections. The second task was a 40-minute open-ended design task where participants were presented with three different startup concepts, selected one, and freely designed a webpage from scratch.
After these tasks concluded, three authors prepared visualizations of the interaction dynamics to facilitate the stimulated recall interview on the second day. The authors first marked the moments when participants sent requests to the agent and when the agent completed execution for each request (Figure 5A). Using these two time points to segment the entire task workflow, the authors visualized the data as shown in Figure 5B-D, distinguishing between periods when the agent was actively working and when it was not. Alongside these visualizations, we prepared segmented video clips of the actual task recordings.
Second day
On the second day, participants reviewed their visualized interaction data and corresponding video clips. For each segmented portion, we inquired about the reasons for their actions in that segment, the purpose behind their actions, why they deemed those actions appropriate, and the types of tasks or collaboration contexts where such actions could be effective. This process took approximately 45 minutes to 1 hour, followed by an in-depth interview lasting approximately 15-20 minutes exploring overall collaboration experiences, the effectiveness of concurrent work, and what implicit user information (e.g., work or collaboration style, or intent) their actions conveyed to the agent.
@@ 4.2.3. Analysis Methods
Qualitative coding of the user’s action patterns.
From the study, we collected a total of 214 segmented processes—that is, 214 turn-level data points from when the agent started a task to when it completed it (Figure 5C). These segments were collected from 10 participants across two tasks, with an average of 21.4 turns per participant in total (min = 12, max = 33). Using this data, three authors first conducted inductive qualitative coding on what actions users performed while the agent was working. When a user performed multiple actions within a single turn, we separated them for individual data. Two authors began by iteratively reviewing the data to consolidate overlapping actions and group similar behaviors into broader categories. Through this iterative process, they organized the data points by merging and splitting them into meaningful groups. Subsequently, a third author joined to examine these groupings, and the three authors collaboratively refined and finalized the category. This iterative refinement continued until all participating authors achieved consensus. As a result, we defined 5 categories (Hands-off, Observational, Terminating, Directive, and Concurrent interaction) and 10 detailed codes (Table 1).
Distribution analysis of user action patterns.
After coding the user action patterns, we returned to the 214 data points and annotated each turn with 5 categories. As shown in Figure 6, each turn was mapped to 1 to 5 categories (Table 1), and we calculated the distribution for each. In one turn, the user could perform multiple action categories. For example, at the start of the agent’s tool call, the user simply watches the agent’s workflow, then leaves to do their own work until the agent finishes. Additionally, we calculated the proportion of turns that included each category across all 214 turns.
Qualitative coding of triggers and enabling factors of user actions.
@@
Next, we conducted qualitative coding to understand when and why users chose different action categories. We analyzed what prompted users to intervene (concurrent, directive, or terminating actions) versus continue observing or delegating, and what contextual factors shaped these choices. Through iterative coding of stimulated recall interview data, we identified two key pattern types: triggers that prompted transitions from observation to intervention, and enabling factors—contextual conditions including mental models, task priorities, and intervention preferences—that determined which specific action users adopted. Specifically, we coded the task and situational characteristics under which each action category occurred. The coding process followed the same methodology used for analyzing user action patterns, with the addition of a fourth author who had not been involved in conducting the study sessions to provide an independent perspective and reduce potential bias.
Iterative decision modeling of human-agent co-creative collaboration.
@@
Building on the identified user action patterns, along with their corresponding triggers and enabling factors, we constructed an initial decision model that describes how user-agent collaboration unfolds from the agent’s task initiation to completion. The model represents transitions between different user actions (e.g., hands-off, observational, directive, and concurrent) and the conditions under which these transitions occur. To validate and refine this model, we iteratively compared it against all 214 turn-level processes. Whenever an observed turn sequence could not be adequately explained by the current model, we revisited the underlying triggers, enabling factors, and transition logic, and revised the model accordingly.
This constant comparative process (Glaser and Strauss, 2017) was conducted by four authors through repeated discussion and refinement, and continued until the decision model consistently accounted for all observed cases from the study without contradiction. The resulting model thus reflects a grounded and data-driven representation of human-agent co-creative workflows, capturing not only interaction patterns but also the mechanisms that give rise to shifts in user involvement over time.

### 07 Apr 2026 v4

@@ 3.2. Study Design
Study 1 investigates two questions: (1) How does process transparency—observing the agent’s step-by-step workflow on the canvas—affect users’ collaboration experience and enable a new type of collaboration? (2) What design considerations are needed for better human-agent collaboration in co-creative workflow? We observed 10 professional UI/UX designers collaborating with our design probe across different task types to examine how process transparency shapes their understanding and collaboration strategies.
@@ 3.2.1. Participants
We recruited 10 UI/UX designers (D1-D10, 3 male, 7 female, ages 24-36, mean=28.30) with at least 1 year of professional design experience. We screened for high Figma proficiency (e.g., daily users) and prior experience with diverse design agents (e.g., Figma Make (Make, ) or Lovart (Lovart, )). Participants were compensated 40,000 KRW for approximately 1.5 hours of participation.
@@ 3.2.2. Procedure
After introducing different types of design agents (e.g., OptiMuse (Zhou et al., 2024) and Figma Make (Make, )) and a brief tutorial on the probe’s three interaction components, participants completed two design tasks of varying scope to observe how collaboration strategies differ across task complexity: (1) a fine-grained task designing a single webpage section (e.g., hero or project gallery) with given materials (15 mins), and (2) a coarse-grained task creating an entire webpage (e.g., startup landing page) across multiple sections (25 mins). We concluded with a 30-minute semi-structured interview exploring how process transparency shaped their collaboration strategies and interaction with the probe.
3.3. Findings
@@ 3.3.1. Process Transparency: Real-Time Understanding and Intervention Opportunities
Observing the agent’s step-by-step execution fundamentally transformed how designers experienced collaboration. All designers reported a strong sense of “working with the agent” rather than delegating to then reviewing it. Unlike traditional agents, where users assign tasks and wait for final outputs, process transparency enables concurrent work: participants understand what the agent is doing, how it will end up, and plan or redirect their own tasks based on this understanding.
This transparency relieved the “loss of agency” problem common in human-agent interaction (Schömbs et al., 2026). Participants contrasted our probe with general design agents, where agents take over the workspace, leaving users waiting passively and fearing disruption if they intervene. With those agents, users must read lengthy reasoning texts to understand their plans or reasoning—a cognitively demanding process. In contrast, seeing intermediate behaviors with output on the canvas made the agent’s intentions intuitively understandable. D4 commented “I don’t need to read paragraph. I just see it creating a button, and I know what’s happening.” This understanding of how the agent works prevented over-dependence on opaque automation, preserving user agency throughout the process.
Most critically, step-by-step observation enabled early evaluation and proactive intervention. Participants could anticipate the agent’s next steps (D1-D10), identify the agent’s misinterpretation on the given task (D1, D3, and D5-D7), and recognize their own unclear instructions (D1-D2, D4, and D8), and assess the agent’s overall workflow direction (D2, D6, and D8-D10)—all before the agent finished. Importantly, this understanding enabled participants to plan their own flexible parallel work: while the agent executed, participants gathered reference images, evaluated overall design directions, or prepared content for subsequent sections, rather than passively waiting. Specifically, D7 stated: “I saw it working on the hero section with a certain direction, so I started looking for references related to that direction for the next section instead of just watching.” This predictability helped participants manage workflows more effectively, deciding when to intervene, to prepare subsequent tasks, and to let the agent continue autonomously. Process transparency thus shifted collaboration from reactive correction of final outputs to proactive coordination during execution.
@@ 3.3.2. Emerging Expectations: Flexible Concurrent Collaboration and Context Awareness
Agent’s process transparency fundamentally changed the users’ expectations: they wanted to engage directly with the agent’s work-in-progress through concurrent intervention rather than waiting for completion. For example, participants wanted to appropriate intermediate outputs for their own designs, assist the agent by working on related elements simultaneously, or demonstrate desired changes through direct editing. Most of the participants (D1-D3 and D6-D10) explained that observing the process made them more aware of how they wanted things done, creating natural moments to intervene and collaborate. D3 noted: “I could see exactly where it needed my input—I wanted to just jump in and show it rather than explain this verbally.”
However, the first probe’s inability to distinguish between user and agent actions in the canvas created conflicts. Because the agent tracked overall canvas changes without separating user modifications, it sometimes reverted or misinterpreted concurrent user work. Participants who experienced this avoided intervening during execution, relegating their collaboration ideas to post-task interviews. Yet even participants who avoided intervention reported that their reluctance to interrupt agents stems from prior experiences with other agent systems (D2, D4-D7, D10)—fearing that mid-execution changes would derail the agent’s workflow. Process transparency challenged this assumption; designers recognized opportune moments for intervention but lacked system support.
Participants envisioned real-time interleaved workflows where users could intervene during the agent’s ongoing execution: appropriating agent-created elements into their own designs while the agent continued working, handing off their work-in-progress for the agent to build upon mid-execution, or collaboratively editing the same artifact simultaneously. Critically, they expected agents “to understand why they intervened”—whether corrections indicated misunderstanding, appropriations showed approval, or demonstrations communicated desired style (D1, D3-D5, D7-D8). This requires the agent’s collaborative context awareness: distinguishing feedback on agent work from independent user work and recognizing when users want the agent to adapt versus continue autonomously. Process transparency thus revealed not just new collaboration opportunities, but the need for context-aware agents that support—rather than conflict with—real-time interleaved human actions.
@@ 4. Study 2: Two-Day Technology Probe Study with Stimulated Recall Interviews
While Study 1 revealed the coordination problem that arises when agents lack collaborative context awareness, it remained unclear what interaction patterns emerge in the collaboration and what enables users to shift between delegation, direction, and concurrent work. To investigate this, we conducted a two-day technology probe study (Hutchinson et al., 2003) with stimulated recall interviews (Dempsey, 2010; Calderhead, 1981) observing experienced creative professionals collaborate with the probe.
4.1. Second Probe System: Cleo
(Collaborative Linked Executive Operator)
@@ 4.1.1. Understanding Collaborative Context for Flexible Human-Agent Co-Creation
Study 1 revealed that process transparency prompts users to interleave their actions with the agent’s ongoing execution. The key challenge is collaborative context awareness—agents must distinguish whether concurrent user actions represent feedback requiring adaptation, independent parallel work to be preserved, or co-editing to be incorporated into ongoing execution.
Therefore, we developed Cleo (Collaborative Linked Executive Operator), which extends the first probe by incorporating collaborative context awareness as its core principle, enabling the agent to recognize and adapt to users’ direct manipulation during execution.
@@ 4.1.2. Cleo Design and Improvements from Design Probe 1
Figure 3. Flexible co-creation scenario with Cleo. (a) The user invokes Cleo by calling its name. (b) While Cleo is acting on the canvas, the user can (c) work concurrently on the same task, or (d) intervene directly with additional instructions at any time.
@@
Interaction Improvements. Building on the first probe system, Cleo supports three additional interaction types (Figure 3). First, users can abort the agent’s operation at any time via the termination button on the plugin. Second, users can work concurrently with the agent on the same elements, with the agent tracking and prioritizing user operations to align its plan accordingly (Figure 3-c). Third, users can provide additional voice instructions at any time by calling the agent’s name while it is working, which the agent incorporates in the next iteration (Figure 3-d).
Technical Improvements. To support these, we replaced the original Summary Module with three new modules (Figure 4). The User Change Detection Module analyzes user actions that occurred while the agent was inactive, maintaining awareness of the overall workflow. The Attribution Change Module distinguishes between agent and user actions during execution, producing separate summaries for each. The Plan Update Module then updates the current plan only when user actions conflict with or indicate new directions from the existing plan, before initiating the next iteration cycle.
Figure 4. Agent structure of Cleo. Three modules have been updated from the first probe pipeline (Appendix A.4): User Change Detection Module, Attribution Change Module, and Plan Update Module.
@@
Figure 5. How Cleo achieves context awareness. (a) The user invokes the agent. (b) It recalls past observations and detects user updates against the current state. (c) It monitors concurrent user actions while executing its plan, prioritizes interventions, and updates its plan as needed.
The figure demonstrates how the system achieves context awareness through three components. In (a), the user invokes the agent by issuing a voice command to add icons to the footer of a portfolio canvas. In (b), the User Change Detection Module compares the agent’s previous observation — a simple footer with email and contact fields — against the current canvas state, detecting that the user has updated the footer content and added contact information. In (c), while the agent is actively creating icons on the canvas, the Attribution Change Module detects that the user concurrently changed the layout alignment to vertical. This intervention triggers the Plan Update Module, allowing the agent to revise its plan accordingly.
@@ 4.2. Study Design
We conducted a two-day exploratory study with Cleo to understand what interaction patterns emerge when users collaborate with a context-aware agent, and what enables or triggers different collaboration modes. By combining interaction logs with stimulated recall interviews, we aimed to capture not just what users did, but why they made those choices in context.
On the first day, participants performed two design tasks (closed- and open-ended) in Figma, with interactions logged and screen-recorded. After the session, three authors visually represented the interaction logs by segmenting the entire task process into agent-activated and agent-deactivated periods, and labeling each segment with timestamped user and agent actions to create an easily reviewable format for the stimulated recall interview ( Appendix B.1-c-d). We gave a one-day gap between the task session and recall interview to avoid participant fatigue and for visualizing logs. On the second day, stimulated recall interviews were conducted where participants reviewed these visualized interaction logs.
@@ 4.2.1. Participants
We recruited 10 participants (4 males and 6 females; M_age = 30.1, SD_age = 3.9) who met two key criteria: (1) at least two years of professional design experience, and (2) active use of VLMs (e.g., ChatGPT (OpenAI, 2026) or Gemini (Google, 2026)) or design agents (e.g., Figma Make (Make, ), Stitch (Stitch, ), or Lovable (Lovable, )) in their creative workflow. Participants were compensated 120,000 KRW for approximately 3 hours of participation.
@@ 4.2.2. Procedure
First day. After an introduction, participants familiarized themselves with Cleo through a warm-up task (30 mins), followed by two main tasks: (1) a closed-ended task (30 mins) where participants designed an empty section of a web portfolio to match existing sections given design requirements and source materials, and (2) a open-ended task (40 mins) where participants selected one of three startup concepts and freely designed a webpage from scratch.
After the tasks, three authors processed the collected interaction logs by marking timestamps of user requests and agent task completion. Using these timestamps, each participant’s entire task process was segmented into agent-activated and agent-deactivated periods, then grouped by each user request. Segmented video clips were also prepared alongside these logs.
After these tasks concluded, three authors prepared visualizations of the interaction dynamics to facilitate the stimulated recall interview on the second day. The authors first marked the moments when participants sent requests to the agent and when the agent completed execution for each request (Appendix B.1-a). Using these two time points to segment the entire task workflow, the authors visualized the data (Appendix B.1-b-d), distinguishing between periods when the agent was actively working and when it was not. Alongside these visualizations, we prepared segmented video clips of the actual task recordings.
Second day. Participants reviewed their visualized interaction logs and corresponding video clips. For each segment, we inquired about the reasons and purpose behind their actions, and the collaboration contexts where such actions could be effective within 1 hour. A 20-minute in-depth interview followed by exploring overall collaboration experiences, the effectiveness of concurrent work, and what implicit information their actions conveyed to the agent.
@@ 4.2.3. Analysis Methods
Qualitative coding of the user’s action patterns. From the study, we collected 214 turn-level data from 10 participants across two tasks, where a turn means a complete agent execution in response to a user request comprising up to 10 iterations (Section 2), with an average of 21.4 turns per participant (min=12, max=33). Three authors conducted open coding (Corbin and Strauss, 1990; Khandkar, 2009) on what actions users performed while the agent was working, separating multiple actions within a single turn for individual coding. Two authors iteratively reviewed the data to consolidate overlapping actions into broader categories, after which a third author joined to refine and finalize until consensus was reached. As a result, we defined 5 categories and 10 codes (Table 1).
Distribution analysis of user action patterns. After coding, we annotated each of the 214 turns with the 5 categories, where each turn could be mapped to 1 to 5 categories (Appendix B.3), and calculated each category’s proportion across all 214 turns.
@@
Qualitative coding of triggers and enabling factors of user actions. Next, we conducted qualitative coding on the stimulated recall interview data to understand when and why users chose different action categories, identifying two key pattern types: triggers that prompted transitions from observation to intervention, and enabling factors—situational conditions including user’s mental models, task priorities, and intervention preferences—that determined which specific action users adopted. Following the same methodology as above, with the addition of a fourth author who had not been involved in conducting the study sessions to provide an independent perspective and reduce potential bias.
@@
Iterative decision modeling of human-agent co-creative collaboration. Building on the identified action patterns, triggers, and enabling factors, we constructed an initial decision model describing how user-agent collaboration unfolds from task initiation to completion, representing transitions between different user actions and the conditions under which these transitions occur. Through a constant comparative process (Glaser and Strauss, 2017), we iteratively validated and refined the model against all 214 turn-level processes, revising it whenever an observed turn sequence could not be adequately explained. This process was conducted by five authors, including an additional author uninvolved in the study sessions to provide an independent perspective, and continued until the model consistently accounted for all cases without contradiction.